find the inverse of f(x)=2^x+1
Answer by fractalier(6550)

Okay, from
f(x)=2^x+1
y = 2^x + 1
x = 2^y + 1
x - 1 = 2^y
log(x-1) = log 2^y
log(x-1) = y log 2
y = log(x-1) / log 2
f^(-1)(x) = log(x-1) / log 2